A 75 L sample of gas is collected at a room temperature of 25.0ºC. What volume will the gas  occupy at 4.66ºC assuming the pressure of the gas remains constant?


1
Expert's answer
2021-02-03T07:19:39-0500

According to Charles' Law at constant pressure;

"\dfrac{V1}{T1}=\dfrac{V2}{T2}= k(Constant)"


V1=75L

T1=25.5+273 = 298K

V2 =?

T2 = 4.66 + 273 = 277.66K

"Thus, V2 = \dfrac{75Lx277.66K}{298K} = 69.88L"